40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9047     Entry time: Wed Aug 21 19:37:25 2013
Author: Masayuki 
Type: Summary 
Category: Green Locking 
Subject: X-arm PDH OLTF measurement 

[Manasa, Masyauki]

Today we measured OLTF of PDH green lock of x-arm again. In the previous measurement the excitation signal was injected at the PDH servo box output(elog 9044), but in this measurement we changed the injection point to the RFPD mixer output (just before the servo input).

We measured the OLTF with the servo gain of 6.5 and source amplitude of 5 mV for all frequency band. The measured UGF was 11 kHz and the phase margin was 48 degree.

Next that measurement, we tried to measure the power spectrum density of the error signal and feedback signal. But the alignment was not so good, so we aligned the green light injection point. Tomorrow we will continue the alignment and will measure the PSD.

attatchment1 - OLTF of PDH green lock with servo gain of 6.5

OLTF_PDH_Glock_6_5_2.png

ELOG V3.1.3-